# GlycanML: A Multi-Task and Multi-Structure Benchmark for Glycan Machine Learning

### v1

@@ 5.3 Benchmark Results on Multi-Task Learning
In Table 4, we report the benchmark results of different MTL methods against single-task learning. We select shallow CNN and RGCN, i.e., best-performing glycan sequence and graph encoders, as the backbone encoder, and all MTL methods are evaluated on each of them. According to benchmark results, we have the following findings:
  • The temperature scaling (TS) approach performs best. On both shallow CNN and RGCN, the TS approach achieves the highest mean accuracy, and it outperforms single-task learning with a clear margin (i.e., 1.63% improvement in mean accuracy) when using RGCN as backbone encoder. Therefore, the TS approach can well balance the learning signals from different glycan taxonomy prediction tasks, leading to stable performance gains.
  • MTL methods are not always beneficial. On shallow CNN, only the naive MTL (N-MTL) and the TS method outperform single-task learning in terms of mean accuracy; on RGCN, only the TS method outperforms single-task learning in terms of mean accuracy. Actually, most MTL methods lead to performance decrease compared to single-task learning. These results suggest the high difficulty of balancing between different glycan taxonomy prediction tasks. More efforts are thus required to boost the MTL performance on the GlycanML-MTL testbed, which we leave as one of our major future works.
  • MTL is more helpful for difficult tasks. From domain-level to species-level classification (i.e., from 4-way classification to 1,737-way classification), the task difficulty monotonically increases. We can observe more benefits of MTL on difficult tasks. For example, on RGCN, the TS method outperforms single-task learning on order-, family-, genus- and species-level classification (i.e., tasks with more categories), while the TS method is inferior on domain-, kingdom-, phylum- and class-level classification (i.e., tasks with fewer categories). Therefore, in MTL, more attention is paid to difficult tasks, leading to better performance on these tasks.
@@ 6 Conclusions and Future Work
In this work, we build a comprehensive benchmark GlycanML for glycan machine learning. It consists of diverse types of glycan understanding tasks, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. In GlycanML, we support two representation methods of glycan structures, i.e., glycan tokenized sequences and glycan planar graphs. Additionally, on eight highly correlated glycan taxonomy prediction tasks, we set up a testbed GlycanML-MTL to compare different multi-task learning (MTL) algorithms. According to the benchmark results, multi-relational graph neural networks (GNNs) show great promise for glycan modeling, and well-designed MTL methods can further boost model performance.
Limitations and future work. The current GlycanML benchmark is limited to modeling glycan sequences and 2D glycan graphs, without benchmark datasets of 3D glycan structures. Therefore, in the future, we will go beyond sequence- and 2D-graph-based datasets and methods to 3D-structure-based ones. Also, we will work along with the community to further promote the efforts of MTL for better glycan understanding.

### v2

@@ 5.3 Benchmark Results on Multi-Task Learning
In Table 4, we report the benchmark results of different MTL methods against single-task learning. We select shallow CNN and RGCN as the backbone encoder, and all MTL methods are evaluated on each of them. According to benchmark results, we have the findings below:
  • The temperature scaling (TS) approach performs best. On both shallow CNN and RGCN, the TS approach achieves the highest mean Macro-F1 score, and it outperforms single-task learning with a clear margin (i.e., 2.75% improvement in mean Macro-F1 score) when using shallow CNN as backbone encoder. Therefore, the TS approach can well balance the learning signals from different glycan taxonomy prediction tasks, leading to stable performance gains.
  • MTL methods are not always beneficial. On shallow CNN, only the TS and UW methods outperform single-task learning in terms of mean Macro-F1 score; on RGCN, only the TS and DTP methods outperform single-task learning in terms of mean Macro-F1 score. Actually, most MTL methods lead to performance decrease compared to single-task learning. These results suggest the high difficulty of balancing between different glycan taxonomy prediction tasks. More efforts are thus required to boost the MTL performance on the GlycanML-MTL testbed, which we leave as one of our major future works.
  • MTL is more helpful for difficult tasks on RGCN. From domain-level to species-level classification (i.e., from 4-way classification to 1,737-way classification), the task difficulty monotonically increases. Under RGCN, we can observe more benefits of MTL on difficult tasks. For example, the DTP method outperforms single-task learning on order-, family, genus- and species-level classification (i.e., tasks with more categories), while the DTP method is inferior on kingdom-, phylum- and class-level classification (i.e., tasks with fewer categories). Therefore, in MTL with RGCN, more attention is paid to difficult tasks, leading to better performance on these tasks.
Table 5: Benchmark results on multi-task learning for glycan function prediction. We report mean (std) for each experiment.
Method Immunogenicity (AUPRC) Glycosylation (Macro-F1) Interaction (Spearman’s $\rho$) Backbone encoder: Shallow CNN Single-Task 0.776(0.027) 0.898(0.009) 0.261(0.008) Multi-Task 0.792(0.028) 0.912(0.011) 0.257(0.002) Backbone encoder: RGCN Single-Task 0.780(0.006) 0.948(0.004) 0.262(0.005) Multi-Task 0.801(0.001) 0.963(0.009) 0.259(0.030)
5.4 The Effect of Multi-Task Learning on Glycan Function Prediction
In this part, we study the effect of MTL on three glycan function prediction tasks, i.e., immunogenicity, glycosylation and interaction prediction. Considering that evolutionarily similar glycans tend to have similar functions, we study how glycan function prediction can be benefited by jointly learning the evolutionary taxonomy of glycans. Specifically, we regard each of three function prediction tasks as the center task and eight taxonomy prediction tasks as auxiliary tasks, and the center task is trained along with eight auxiliary tasks to perform MTL (the losses of center and auxiliary tasks are summed up for model training).
In Table 5, we compare this multi-task method with the single-task baseline. Under both Shallow CNN and RGCN encoders, MTL performs better on immunogenicity and glycosylation type prediction, while single-task learning performs better on interaction prediction. These results show the potential of MTL on boosting glycan function prediction, and also inspire future research efforts to further improve its effectiveness.
@@ 6 Conclusions and Future Work
In this work, we build a comprehensive benchmark GlycanML for glycan machine learning. It consists of diverse types of glycan understanding tasks, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. In GlycanML, we support two representation methods of glycans, i.e., glycan tokenized sequences and glycan planar graphs. Additionally, on eight highly correlated glycan taxonomy prediction tasks, we set up a testbed GlycanML-MTL to compare different multi-task learning (MTL) algorithms. Also, we study how taxonomy prediction can boost other three function prediction tasks by MTL. According to the benchmark results, multi-relational graph neural networks (GNNs) show great promise for glycan modeling, and well-designed MTL methods can further boost model performance.
Limitations and future work. The current GlycanML benchmark is limited to modeling glycan sequences and 2D glycan graphs, without benchmark datasets of 3D glycan structures. Therefore, in the future, we will go beyond sequence- and 2D-graph-based datasets and methods to 3D-structure-based ones [5, 16]. Also, we will work along with the community to further promote the efforts of MTL for better glycan understanding.